# Leveraging Joint Predictive Embedding and Bayesian Inference in Graph Self Supervised Learning

### v1

@@ 4.2.1 Performance on Node Classification
GCN $81.5\pm 1.3$ $\underline{71.9\pm 1.9}$ $77.8\pm 2.9$ $91.2\pm 1.2$ $\underline{82.6\pm 2.4}$ $91.1\pm 0.5$
GAT $\underline{81.8\pm 0.3}$ $71.4\pm 1.8$ $78.7\pm 2.3$ $85.7\pm 20.3$ $78.0\pm 19.0$ $90.5\pm 0.6$
Simplified GCN $81.5\pm 0.2$ $73.1\pm 0.1$ $\underline{79.7\pm 0.4}$ $88.3\pm 1.1$ N/A $\underline{91.5\pm 0.3}$
GraphSage Mean $79.2\pm 7.7$ $71.6\pm 1.9$ $77.4\pm 2.2$ $\underline{91.4\pm 1.3}$ $82.4\pm 1.8$ $91.3\pm 2.8$
@@
Ours $\mathbf{89.8\pm 0.9}$ $\mathbf{77.0\pm 0.9}$ $\mathbf{85.7\pm 1.0}$ $\mathbf{94.5\pm 0.5}$ $\mathbf{88.0\pm 0.6}$ $\mathbf{93.6\pm 0.4}$
Table 1: Supervised Node Classification. The values indicate the accuracy achieved on node classification by several methods. A higher value indicates better performance. The best score is marked in bold and the second best score is underlined. N/A indicates that the score was not reported for a particular dataset by the original authors.
Method Cora Citeseer Pubmed Amazon P Amazon Comp. Coauthor CS
DGI $81.7\pm 0.6$ $71.5\pm 0.7$ $77.3\pm 0.6$ $83.1\pm 0.3$ $83.6\pm 0.2$ $90.0\pm 0.3$
MVGRL $82.9\pm 0.7$ $72.6\pm 0.7$ $79.4\pm 0.3$ $87.3\pm 0.1$ $82.8\pm 0.1$ $91.3\pm 0.1$
GRACE $80.0\pm 0.4$ $71.7\pm 0.6$ $79.5\pm 1.1$ $81.8\pm 0.8$ $89.5\pm 0.3$ $71.1\pm 0.2$
CCA-SSG $84.2\pm 0.4$ $73.1\pm 0.3$ $81.6\pm 0.4$ $93.1\pm 0.1$ $88.7\pm 0.3$ $\underline{93.3\pm 0.2}$
SUGRL $83.4\pm 0.5$ $73.0\pm 0.4$ $81.9\pm 0.3$ $\underline{93.2\pm 0.4}$ $\underline{88.9\pm 0.2}$ $92.2\pm 0.5$
S3-CL $\underline{84.5\pm 0.4}$ $\underline{74.6\pm 0.4}$ $80.8\pm 0.3$ $89.0\pm 0.5$ N/A $93.1\pm 0.4$
GraphMAE $84.2\pm 0.4$ $73.1\pm 0.4$ $83.9\pm 0.3$ $90.7\pm 0.4$ $79.4\pm 0.5$ $93.1\pm 0.1$
GMI $82.7\pm 0.2$ $73.3\pm 0.3$ $77.3\pm 0.6$ $83.1\pm 0.3$ N/A $91.0\pm 0.0$
BGRL $83.8\pm 1.6$ $72.3\pm 0,9$ $\mathbf{86.0\pm 0.3}$ $93.2\pm 0.3$ $\mathbf{90.3\pm 0.2}$ $93.3\pm 0.1$
Ours $\mathbf{89.8\pm 0.9}$ $\mathbf{77.0\pm 0.9}$ $\underline{85.7\pm 1.0}$ $\mathbf{94.5\pm 0.5}$ $88.0\pm 0.6$ $\mathbf{93.6\pm 0.4}$
@@
Table 2: Self Supervised followed by Fine-Tuning. The values indicate the accuracy achieved on node classification by several SSL methods. A higher value indicates better performance. The best score is marked in bold and the second best score is underlined. N/A indicates that the score was not reported for a particular dataset by the original authors.
Method Photos Computers CS
BGRL $93.2\pm 0.3$ $90.3\pm 0.2$ $93.3\pm 0.1$
ParetoGNN $93.8\pm 0.3$ $\mathbf{90.7\pm 0.2}$ $92.2\pm 0.1$
Ours $\mathbf{94.5\pm 0.5}$ $88.0\pm 0.6$ $\mathbf{93.6\pm 0.4}$
Table 3: Comparision with BGRL and ParetoGNN on datasets introduced in [19]. The best score for each dataset is marked in bold.
4.2.2 Study on GMM Cluster Optimization
@@
In this section, we evaluate the significance of the constraint in our loss function. Our approach combines learning sub-graph embeddings with optimizing node embeddings using pseudo-labels derived from a Gaussian Mixture Model (GMM). These pseudo-labels, converted to normalized scores, serve as a constraint in the original loss function for subgraph embeddings. Table 4 illustrates the performance of our method with and without this constraint, demonstrating a substantial improvement when the GMM-derived pseudo-label scores are incorporated. This finding is particularly noteworthy as it indicates that computationally expensive processes like negative sampling, commonly used in graph contrastive learning, may not be essential for optimizing node embeddings. Our approach thus offers a more efficient alternative while maintaining, and even enhancing, performance in graph representation learning tasks.
Method Cora Citeseer Pubmed Amazon P Amazon Comp. Coauthor CS
Without Bayesian Inference 89.0 74.1 84.2 93.4 85.0 93.2
With Bayesian Inference $\mathbf{89.8}$ $\mathbf{77.0}$ $\mathbf{85.7}$ $\mathbf{94.5}$ $\mathbf{88.0}$ $\mathbf{94.5}$
@@
Table 4: Ablation Study for Loss Function.The best score for each dataset is marked in bold.
4.2.3 Study on Momentum Parameter $m$
@@
Figure 2 represents the change in performance with respect to the momentum parameter on the citeseer dataset. The ideal values for $m$ may range between 0.8 and 0.9. As $m\rightarrow 1$, the weights of the target encoder update by very small steps, thereby slowing down learning. This is further indicated by the drop in performance. For all our experiments, we use $m=0.9$
4.2.4 Spread of Embeddings
Figure 3 represents the spread of node embeddings across all 6 tested datasets. As mentioned earlier, the node embeddings are bounded between -1 and 1 since the final layer activation function is Tanh. The plots indicate that the node embeddings are evenly distributed with a high dispersion indicating a higher variance, implying that the embeddings span a wide range of values. This is a key indicator that our proposed method effectively prevents node representations from clustering too closely, ensuring that each node maintains a distinct position within the embedding space. This controlled spread minimizes overlap between node embeddings, thereby enhancing the uniqueness and robustness of the overall representation.
Figure 3: Box plot representing the learned embeddings on all tested datasets
4.2.5 Study on Performance against Test Time Node Feature Distortion
We evaluate the performance of our method on node classification by augmenting the test splits. We sample $p$ percentage test nodes randomly and replace their node features with points sampled from a standard Gaussian distribution. We evaluate the performance of the distorted test nodes on Amazon Computers, Amazon Photos and Coauthor CS by varying $p$ between 10% to 40% and report their performance differences in table 5. In this study, we only use Co-author and Amazon networks since citation networks tend to be unstable on evaluation as indicated in [19]. Despite no further fine-tuning on distorted node features, the average performance drop on Amazon Photos and Computers is only 2.96% and 4.01% respectively. This indicates the ability of the joint framework to generalize on noise-augmented views as well. According to our hypothesis, we believe this is due to the nature of joint predictive embedding learning where a higher variation in context and target embeddings is required for better generalization.
Ratio of Nodes Distorted Photos Computers Coauthor CS
0.10 -3.3% -3.8% -3.3%
0.15 -0.8% -3.4% -1.2%
0.20 -3.7% -3.3% -3.3%
0.25 -2.9% -4.6% -7.7%
0.30 -3.9% -4.3% -5.9%
0.35 -2.4% -4.0% -6.6%
0.40 -3.7% -4.7% -9.6%
Table 5: Performance on distorting node features at test time. Values represent the percentage decrease in the model’s performance compared to its original score. A lower value indicates better performance.
5 Conclusion

### v2

@@ IV-B1 Performance on Node Classification on small and large graphs
GCN $81.5\pm 1.3$ $71.9\pm 1.9$ $77.8\pm 2.9$ $91.2\pm 1.2$ $82.6\pm 2.4$ $91.1\pm 0.5$
GAT $81.8\pm 0.3$ $71.4\pm 1.8$ $78.7\pm 2.3$ $85.7\pm 20.3$ $78.0\pm 19.0$ $90.5\pm 0.6$
Simplified GCN $81.5\pm 0.2$ $73.1\pm 0.1$ $79.7\pm 0.4$ $88.3\pm 1.1$ N/A $91.5\pm 0.3$
GraphSage Mean $79.2\pm 7.7$ $71.6\pm 1.9$ $77.4\pm 2.2$ $91.4\pm 1.3$ $82.4\pm 1.8$ $91.3\pm 2.8$
@@
Ours 89.8 ± 0.9 77.0 ± 0.9 85.7 ± 1.0 94.5 ± 0.5 88.0 ± 0.6 94.5 ± 0.4
@@
TABLE III: Self Supervised Pre-Training followed by classification on Planetoid Datasets. The values indicate the accuracy achieved on node classification by several SSL methods. A higher value indicates better performance. The best score is marked in bold and the second-best score is underlined.
Method Cora Citeseer Pubmed
DGI $81.7\pm 0.6$ $71.5\pm 0.7$ $77.3\pm 0.6$
MVGRL $82.9\pm 0.7$ $72.6\pm 0.7$ $79.4\pm 0.3$
GRACE $80.0\pm 0.4$ $71.7\pm 0.6$ $79.5\pm 1.1$
CCA-SSG $84.2\pm 0.4$ $73.1\pm 0.3$ $81.6\pm 0.4$
SUGRL $83.4\pm 0.5$ $73.0\pm 0.4$ $81.9\pm 0.3$
S3-CL $\underline{84.5\pm 0.4}$ $\underline{74.6\pm 0.4}$ $80.8\pm 0.3$
GraphMAE $84.2\pm 0.4$ $73.1\pm 0.4$ $83.9\pm 0.3$
GMI $82.7\pm 0.2$ $73.3\pm 0.3$ $77.3\pm 0.6$
BGRL $83.8\pm 1.6$ $72.3\pm 0.9$ $\mathbf{86.0\pm 0.3}$
Ours $\mathbf{89.8\pm 0.9}$ $\mathbf{77.0\pm 0.9}$ $\underline{85.7\pm 1.0}$
TABLE IV: Self Supervised Pre-Training followed by classification on datasets mentioned in [21] and WikiCS. The best score for each dataset is marked in bold and the second-best score is underlined.
Method Photos Computers Coauthor CS WikiCS
DGI $91.6\pm 0.2$ $83.9\pm 0.5$ $92.1\pm 0.6$ $75.3\pm 0.1$
GRACE $92.1\pm 0.5$ $86.7\pm 0.8$ $93.2\pm 0.4$ $77.5\pm 0.6$
BGRL $93.2\pm 0.3$ $90.3\pm 0.2$ $93.3\pm 0.1$ $80.0\pm 0.1$
ParetoGNN $93.8\pm 0.3$ $\mathbf{90.7\pm 0.2}$ $92.2\pm 0.1$ $\mathbf{82.9\pm 0.1}$
MVGRL $93.2\pm 0.3$ $87.5\pm 0.1$ $92.1\pm 0.1$ $77.5\pm 0.1$
Random Weights $92.1\pm 0.5$ $86.5\pm 0.4$ $91.6\pm 0.3$ $78.9\pm 0.6$
Ours $\mathbf{94.5\pm 0.5}$ $88.0\pm 0.6$ $\mathbf{94.5\pm 0.4}$ $82.4\pm 1.0$
IV-B2 Study on GMM Cluster Optimization
@@
In this section, we evaluate the significance of the constraint in our loss function. Our approach combines learning sub-graph embeddings with optimizing node embeddings using pseudo-labels derived from a Gaussian Mixture Model (GMM). These pseudo-labels, converted to normalized scores, serve as a constraint in the original loss function for subgraph embeddings. Table V illustrates the performance of our method with and without this constraint, demonstrating a substantial improvement when the GMM-derived pseudo-label scores are incorporated. This finding is particularly noteworthy as it indicates that computationally expensive processes like negative sampling, commonly used in graph contrastive learning, may not be essential for optimizing node embeddings. Our approach thus offers a more efficient alternative while maintaining, and even enhancing, performance in graph representation learning tasks.
@@
TABLE V: Ablation Study for Loss Function. The best score for each dataset is marked in bold.
Method Cora Citeseer Pubmed Photos Computers Coauthor CS WikiCS
Without Bayesian Inference 89.0 74.1 84.2 93.4 85.0 93.2 81.5
With Bayesian Inference 89.8 77.0 85.7 94.5 88.0 94.5 82.4
IV-B3 Study on Momentum Parameter $m$
Figure 2 represents the change in performance with respect to the momentum parameter on the citeseer dataset. The ideal values for $m$ may range between 0.8 and 0.9. As $m\rightarrow 1$, the weights of the target encoder update by very small steps, thereby slowing down learning. This is further indicated by the drop in performance. For all our experiments, we use $m=0.9$.
@@ IV-B3 Study on Momentum Parameter $m$
IV-B4 Efficiency Analysis for Scalability
We evaluate the model’s efficiency in terms of the parameter size and memory consumption. We show that our model is among the most efficient of all baselines while outperforming most baselines. We compare the efficiency of our proposed model against reported baselines in table VI.
TABLE VI: Efficiency Comparison among Baselines on Cora, Pubmed, and Citeseer. Lower the number of parameters, more efficient is the model
Methods Cora Citeseer Pubmed
Memory (GB) Params Memory (GB) Params Memory (GB) Params
DGI 3.73 $7.3\times 10^{5}$ 3.85 $1.9\times 10^{6}$ 3.66 $2.6\times 10^{5}$
GMI 4.06 $9.9\times 10^{5}$ 4.20 $2.2\times 10^{6}$ 3.93 $5.2\times 10^{5}$
MVGRL 2.25 $9.9\times 10^{5}$ 2.55 $2.2\times 10^{6}$ 2.37 $5.2\times 10^{5}$
SUGRL 1.57 $9.7\times 10^{5}$ 1.71 $2.6\times 10^{6}$ 1.70 $3.9\times 10^{5}$
S3-CL 1.37 $7.3\times 10^{5}$ 1.62 $1.9\times 10^{6}$ 1.54 $2.6\times 10^{5}$
Ours 0.70 $3.5\times 10^{5}$ 0.80 $6.4\times 10^{5}$ 1.90 $2.3\times 10^{5}$
IV-B5 Study on Performance against Test Time Node Feature Distortion
We evaluate the performance of our method on node classification by augmenting the test splits. We sample $p$ percentage test nodes randomly and replace their node features with points sampled from a standard Gaussian distribution. We evaluate the performance of the distorted test nodes on Amazon Computers, WikiCS, Amazon Photos and Coauthor CS by varying $p$ between 10% and 40% and report their performance differences in table VII. In this study, we only use WikiCS, Co-author and Amazon networks since Planetoid datasets tend to be unstable on evaluation as indicated in [21]. Despite no further fine-tuning on distorted node features, the average performance drop on Amazon Photos and Computers is only 2.96% and 4.01% respectively. This indicates the ability of the joint framework to generalize on noise-augmented views as well. According to our hypothesis, we believe this is due to the nature of joint predictive embedding learning where a higher variation in context and target embeddings is required for better generalization.
TABLE VII: Performance on distorting node features at test time. Values represent the percentage decrease in the model’s performance compared to its original score. A lower value indicates better performance.
Ratio Photos Computers Coauthor CS WikiCS
0.10 -3.3% -3.8% -3.3% -8.9%
0.15 -0.8% -3.4% -1.2% -9.7%
0.20 -3.7% -3.3% -3.3% -13.6%
0.25 -2.9% -4.6% -7.7% -12.6%
0.30 -3.9% -4.3% -5.9% -13.2%
0.35 -2.4% -4.0% -6.6% -12.3%
0.40 -3.7% -4.7% -9.6% -15.2%
V Conclusion